# Mamba Meets Financial Markets: A Graph-Mamba Approach for Stock Price Prediction

### v1

@@ 4 Performance Evaluation
5 Conclusion
The complex nature of stock markets makes predicting stock returns challenging. While transformer-based models have shown promising results, they require high computational resources. We introduce SAMBA, a computationally more efficient but still accurate solution for stock prediction. SAMBA leverages the selective scan nature of the Mamba architecture to capture long-term patterns and uses adaptive graph convolution to infer the relationships between daily stock features. Our experiments show that SAMBA significantly outperforms state-of-the-art baselines in prediction accuracy with low computational complexity, making it a valuable tool for real-time trading and financial analysis. For future work, we will investigate multi-step joint stock price prediction by considering stock correlations.

### v2

@@ IV Performance Evaluation
V Conclusion
The complex nature of stock markets makes predicting stock returns challenging. While transformer-based models have shown promising results, they require high computational resources. We introduce SAMBA, a computationally more efficient but still accurate solution for stock return prediction. SAMBA leverages the selective scan nature of the Mamba architecture to capture long-term patterns and uses adaptive graph convolution to infer the relationships between daily stock features. Our experiments show that SAMBA significantly outperforms several state-of-the-art baselines in prediction performance with low computational complexity, making it a valuable tool for real-time trading and financial analysis. For future work, we will investigate multi-step joint stock price prediction by considering correlations of stock markets.